X^2y"+x(x-1) y'+(1-x)y=0 solve by series solution
1
Expert's answer
2020-04-22T19:41:06-0400

Problem:

"x^2\dfrac{\partial^2 y}{\partial x^2}+x(x-1)\dfrac{\partial y}{\partial x}+(1-x)y=0"  (1)

Let's image solution of the task in the next view:

"y=x^\sigma(a_{0}+a_{1}x+a_{2}x^2+...a_{k}x^k+...)" (2)

Now subsitute equation (2) into equation (1):

"x^2
(\sigma\times(\sigma-1)\times a_{0}x^{\sigma-2} + 
(\sigma+1)\times\sigma\times a_{1}x^{sigma-1} +
(\sigma+2)\times(\sigma+1)\times a_{2} x^{\sigma}+ ...) +
x(x-1)
(\sigma\times a_{0}x^{\sigma-1} +
(\sigma+1)\times a_{1}x^{\sigma} +
(\sigma+2)\times a_{2}x^{\sigma+1}+....) + 
(1-x)
(a_{0}x^{\sigma} + a_{1}x^{\sigma+1} + a_{2}x^{\sigma+2}+...)=0"

Or:

"(a_{0}x^{\sigma}\times\sigma\times(\sigma-1) +a_{1}x^{\sigma+1}\times\sigma\times(\sigma+1)+a_2x^{\sigma+2}\times(\sigma+1)\times(\sigma+2)+...) +
(a_0x^{\sigma+1}\times\sigma+a_1x^{\sigma+2}\times(\sigma+1)+a_2x^{\sigma+3}\times(\sigma+2)+...) - 
(a_0x^\sigma\times\sigma+a_1x^{\sigma+1}\times{\sigma+1}+a_2x^{\sigma+2}\times(\sigma+2)+...) +
(a_0x^{\sigma}+a_1x^{\sigma+1}+a_2x^{\sigma+2}+...)-(a_0x^{\sigma+1}+a_1x^{\sigma+2}+a_2x^{\sigma+3}+...)=0"

Now equate all coefficient near "x^{\sigma}, x^{\sigma+1}, x^{\sigma+2},..."  to zero:

"\begin{cases}
a_0\times\sigma\times(\sigma-1)-a_0\times\sigma+a_0=0\\
a_1\times\sigma\times(\sigma+1)+a_0\times\sigma-a_1\times(\sigma+1)+a_1-a_0=0\\
a_2\times(\sigma+1)\times(\sigma+2)+a_1\times(\sigma+1)-a_2\times(\sigma+2)+a_2-a_1=0\\
...
\end{cases}"

From the first equation (we take that "a_0\not=0" )we get: "(\sigma-1)^2=0=>\sigma=1"

Let's substitute this value to our system (3). Now we have:

"\begin{cases}
a_0\not=0 \\
a_1=0 \\
a_2=0\\
...
\end{cases}"

So we have that all coefficient "a_k" (except "a_0" ) are equal zero. That means that:

"y=a_0x^\sigma"

Also we know that "\sigma=1". So solution of this problem is:

"y=a_0x",

where "a_0" is a different constant.